# Context Embeddings for Efficient Answer Generation in RAG

### v1

@@ 4.5. Baselines with Context Compression
  • ICAE.  (Ge et al., 2024): We use the Mistral-7B-Instruct-v0.2 LoRa-checkpoint [^9] https://huggingface.co/sggetao/icae which uses the same base LLM as ours and is therefore directly comparable. ICAE is fine-tuned to compress a single long context, however, in our work we use multiple contexts. To alleviate this we concatenate the top five retrieved documents together as the context input for the model and truncate as the maximum length of 512 tokens. Note the model has a maximum output length of 128 compressed tokens, which approximately indicates a compression rate of 4 from its original concatenated context input.
  • xRAG. We utilize the xRAG-7b [^10] https://huggingface.co/Hannibal046/xrag-7b, and 8x7B mixture-of-experts model  [^11] https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1 alongside their strongest SFR compressor. The base model is again the same as ours, to ensure comparability. As their model is limited to compressing only a single context into a single compressed representation, we use the top retrieved document for the xRAG setting. [^12] We also tested compressing five documents together, which yielded lower effectiveness. We apply their predefined stopping criteria for answer generation, which aims at cutting the verbose nature of a untuned decoder LLM.
@@ 5.2. COCOM-light
To alleviate the dimensional mismatch between the bert-based compressor and the - typically larger - LLM, we learn a linear projection layer $\bm{W}^{\gamma\times d}$, where $\gamma$ is the compression rate and $d$ the hidden dimension of the LLM. To obtain a set of Context Embeddings we leverage the last hidden representation of each input token. We simply split the hidden representations into blocks of length $\gamma$ and project each block into a single Context Embedding. This way, we learn a block-wise aggregation of the input representations that depending on the input length, and the compression rate $\gamma$ yields a different number of Context Embeddings per input. Note that a similar approach is applied in xRAG, where a projection layer is used on the embedding vector to resolve the dimensional mismatch. However, we argue that compressing using a single vector embedding could significantly restrict the compression quality, especially when using lightweight encoder models such as BERT. This restriction can result in much lower effectiveness compared to using a larger embedding model (Cheng et al., 2024).
@@ 5.3. Computational Efficiency
Table 3. Decoding efficiency in generation Time, GPU Memory, and number of operations (GFLOPs) for COCOM (-light) on dataset NQ. $\xi$ the compression rate. Increasing times is compared against RAG(no compr.).
@@ 5.4. Ablations
Table 5. Impact of the number of provided contexts on COCOM  measured in EM on datasets NQ and ASQA.
@@ 6.1. Context compression
Our results in Table 7 indicate that both models effectively learn the auto-encoding task at lower compression rates ($\xi$=4, 16), but struggle to recover the input when the context is compressed into fewer embeddings ($\xi$=128), with this issue being more pronounced for the lightweight compression module.
We identify two possible explanations: First, compressing longer contexts is inherently more challenging and might require additional objectives. Second, decompressing information from a smaller set of Context Embeddings may be more difficult due to the sequential decoding nature of decoder-only models. Introducing additional pause tokens (Goyal et al., 2024) could help alleviate this issue, providing the model with a means to hierarchically decompress information, drawing on ideas from Chain-of-Thought prompting (Wei et al., 2022). We also experimented with pre-training on more samples but found that this did not improve downstream performance. Regarding the second pre-training task, it is noteworthy that COCOM-light outperforms its larger counterpart in language modeling from Context Embeddings. This analysis shows compressing and re-constructing longer texts is challenging and needs further investigation.

### v2

@@ 4.5. Baselines with Context Compression
  • ICAE.  (Ge et al., 2024): We use the Mistral-7B-Instruct-v0.2 LoRa-checkpoint [^9] https://huggingface.co/sggetao/icae which uses the same base LLM as ours and is therefore directly comparable. ICAE is fine-tuned to compress a single long context, however, in our work we use multiple contexts. To alleviate this we concatenate the top five retrieved contexts together as the context input for the model and truncate as the maximum length of 512 tokens. Note the model has a maximum output length of 128 compressed tokens, which approximately indicates a compression rate of 4 from its original concatenated context input.
  • xRAG. We utilize the xRAG-7b [^10] https://huggingface.co/Hannibal046/xrag-7b, and 8x7B mixture-of-experts model  [^11] https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1 alongside their strongest SFR compressor. The base model is again the same as ours, to ensure comparability. As their model is limited to compressing only a single context into a single compressed representation, we use the top retrieved context for the xRAG setting. [^12] We also tested compressing five contexts together, which yielded lower effectiveness. Again, their compressed representations stem from their dense retriever and are only adapted to the task through a simple linear projection which might limit the model its ability to compress contexts effectively. We apply their predefined stopping criteria for answer generation, which aims at cutting the verbose nature of a untuned decoder LLM.
@@ 5.2. COCOM-light
To alleviate the dimensional mismatch between the bert-based compressor and the - typically larger - LLM, we learn a linear projection layer $\bm{W}^{\gamma b\times d}$, where $\gamma$ is the compression rate, $b$ is the hidden dimension of BERT, and $d$ the hidden dimension of the LLM. To obtain a set of Context Embeddings we leverage the last hidden representation of each input token. We simply split the hidden representations into blocks of length $\gamma$ and project each block into a single Context Embedding. This way, we learn a block-wise aggregation of the input representations that depending on the input length, and the compression rate $\gamma$ yields a different number of Context Embeddings per input. Note that a similar approach is applied in xRAG, where a projection layer is used on the embedding vector to resolve the dimensional mismatch. However, we argue that compressing using a single vector embedding could significantly restrict the compression quality, especially when using lightweight encoder models such as BERT. This restriction can result in much lower effectiveness compared to using a larger embedding model (Cheng et al., 2024).
@@ 5.3. Computational Efficiency
Table 3. Decoding efficiency in generation Time, GPU Memory, and number of operations (GFLOPs) for COCOM (-light) on dataset NQ. $\xi$ the compression rate. Efficiency speedup compared against RAG (no compression) is indicated in brackets.
@@ 5.4. Ablations
Table 5. Impact of the number of provided contexts (k) on COCOM  measured in EM on datasets NQ and ASQA.
@@ 6.1. Context compression
Table 7 showcases the results of these evaluations. Both the full and lightweight models effectively master the auto-encoding task at lower compression factors ($\xi=4,16$). However, they exhibit significant difficulties in reconstructing the input when the compression ratio increases ($\xi=128$). This problem is notably more pronounced in our decoder-based compression model (COCOM).
We identify two possible explanations: First, compressing longer contexts into fewer embeddings inherently presents a challenge due to the inevitable information loss at higher compression rates. Second, the dimension of linear projection layers in the COCOM-light model is dependent on the compression rate; thus, a higher compression rate results in an increased parameter count within its linear layer to manage context compression. In contrast, the COCOM model employs lora tuning, where the size of the components is not dependent on the compression rate. This fundamental difference in handling compression may explain why the COCOM-light model could potentially achieve higher effectiveness under conditions of high compression, due to its higher parameter count. In terms of the second pre-training task, our results indicate that COCOM consistently outperforms COCOM-light, this finding also correlates to the final effectiveness of question-answering tasks, as indicated in table 2.